OpenStudy (anonymous):

Solve for m. -3+m/9=10 -30 63 87 93

OpenStudy (dbzfan836):

93 is the answer

OpenStudy (dbzfan836):

use the properties of equality m+-3/9=10 m+-3/9*9=10*9 m+-3+3=90+3 m=93

OpenStudy (dbzfan836):

bye